Main page‎ > ‎Observing Runs‎ > ‎Night Reports‎ > ‎

Night Report 2014-09-03

 Observers: Leonardo Tartaglia, Nancy Elias-Rosa
 Support: Enrico Cappellaro
 TAT on duty: Maria Teresa Botticella, Andrea Pastorello

--------------------------------------------------------------

 WEATHER CONDITIONS:
 start: wind 8.9 m/s; temp. 7.9 C; humid. 40.0 % - Clear
 middle: wind 14.0 m/s; temp. 7.3 C; humid. 33.0 % - Clear
 end: wind 15.0 m/s; temp. 5.1 C; humid. 32.4 % - Clear

--------------------------------------------------------------

 INSTRUMENTS: SOFI + EFOSC2

--------------------------------------------------------------

 CALIBRATIONS

 Afternoon:
 -- SOFI Dome JHK flats, and BG/RG flats 

--------------------------------------------------------------

 OBSERVATIONS

 Target                                             Type                     UT-Start              Filter/Grism         Slit (")         Seeing (")         Notes

 ** SOFI **
 
 CSS140424-133007-212728          Follow-up             23:20                  Ks                         --             2.0                   (1)
 SN2009ip                                        Follow-up             00:52                   JHKs                     --             1.1                  
 Feige110                                         Standard               02:33                  BG+RG                  1.0            0.9        
 ASASSN14gm                                   Follow-up            02:50                   BG+RG                 1.0            0.9                   (2)
 NGC7720OT                                    Follow-up             05:48                  Ks                         --             1.3                   (3)
 

 ** EFOSC2 **

 OGLE-2014-SN-047                        Follow-up             07:18                  Gr#13                  1.0            1.28                          
 LSQ14eez                                        Follow-up             08:25                  Gr#13                   1.0            1.00                      
 ASASSN14gm                                   Follow-up             09:38                  Gr#13                  1.0            1.26             
 Vma2                                               Standard               10:01                  Gr#11+Gr#13      1.0            1.24

 (1) SN not clearly visible in the combined images                  
 (2) OB repeated, in the second we increased the exptime of the RG spectrum
 (3) OB stopped due to strong wind


 06:20 UT: observations suspended for strong wind from NE. 
 07:03 UT: due to pointing limitations we shifted to EFOSC2

Comments